# Exploring the Robustness and Transferability of Patch-Based Adversarial Attacks in Quantized Neural Networks

### v1

@@ Abstract
Quantized neural networks (QNNs) are increasingly used for efficient deployment of deep learning models on resource-constrained platforms, such as mobile devices and edge computing systems. While quantization reduces model size and computational demands, its impact on adversarial robustness—especially against patch-based attacks—remains inadequately addressed. Patch-based attacks, characterized by localized, high-visibility perturbations, pose significant security risks due to their transferability and resilience. In this study, we systematically evaluate the vulnerability of QNNs to patch-based adversarial attacks across various quantization levels and architectures, focusing on factors that contribute to the robustness of these attacks. Through experiments analyzing feature representations, quantization strength, gradient alignment, and spatial sensitivity, we find that patch attacks consistently achieve high success rates across bitwidths and architectures, demonstrating significant transferability even in heavily quantized models. Contrary to the expectation that quantization might enhance adversarial defenses, our results show that QNNs remain highly susceptible to patch attacks due to the persistence of distinct, localized features within quantized representations. These findings underscore the need for quantization-aware defenses that address the specific challenges posed by patch-based attacks. Our work contributes to a deeper understanding of adversarial robustness in QNNs and aims to guide future research in developing secure, quantization-compatible defenses for real-world applications.
@@ 1 Introduction
As deep learning models are increasingly deployed in resource-constrained environments like mobile devices and edge computing platforms, quantization has emerged as a critical method for reducing model size and computational load. By lowering the bitwidth of weights and activations, quantization enables models to operate more efficiently, preserving accuracy while significantly reducing resource demands [28]. Quantized neural networks (QNNs) are now central to applications ranging from autonomous systems [25, 18] to healthcare [35] and Internet of Things (IoT) devices [32, 12, 16]. However, as QNNs gain prominence, ensuring their robustness against adversarial attacks—long-standing vulnerabilities in deep learning—remains an essential concern for secure deployment.
Among adversarial threats, patch-based attacks pose a unique and formidable challenge [9]. Unlike conventional adversarial perturbations that rely on subtle, dispersed changes, patch attacks introduce highly visible, localized perturbations that disrupt model predictions by “hijacking” the model’s focus. These attacks are especially problematic because of their transferability and resilience across diverse configurations, making them highly effective in real-world, physically realizable scenarios. The distinct nature of patch-based attacks raises pressing questions about the robustness of QNNs against these localized, high-intensity adversarial patterns.
Quantization is often viewed as a potential defense against adversarial attacks, with evidence suggesting that reduced precision may mitigate some types of adversarial perturbations [3, 30, 7]. However, the effectiveness of quantization in countering patch-based attacks remains largely untested. Given that patch attacks rely on distinct, highly visible features to exploit model vulnerabilities, it is reasonable to suspect that they may retain their efficacy even in quantized models. Most existing studies on adversarial robustness in QNNs have focused on dispersed or pixel-level attacks, overlooking the unique challenges presented by patch-based attacks.
@@
  • We systematically evaluate how patch-based attacks affect feature representations across different quantization levels. Our findings reveal that quantization fails to effectively disrupt the adversarial signal of patches, as their high-visibility features persist across bit widths. This persistence underscores a critical vulnerability in QNNs.
@@
  • Based on our findings, we propose key strategies to enhance the robustness of QNNs against patch-based adversarial attacks.
@@ 2 Related Work
QNNs are widely adopted to reduce model size and computational costs, making them suitable for edge and mobile device applications [28]. Quantization-Aware Training (QAT) [13, 29, 36, 23] and Post-Training Quantization (PTQ) [27, 24, 14, 33] are two prominent approaches for implementing low-bitwidth models, each introducing varying degrees of robustness against adversarial attacks [3, 30, 7]. Recent studies indicate that while quantization may disrupt the gradient landscape, creating potential resistance to some pixel-level perturbations [8, 26, 5], its effectiveness against more structured adversarial patterns remains uncertain [35]. Research on QNNs has largely centered on dispersed, pixel-level adversarial attacks [21, 34], leaving a gap in understanding how QNNs respond to localized, high-visibility attacks like patch-based adversarial patterns.
@@ 3 Methodology
In this study, we systematically investigate the robustness and transferability of patch-based adversarial attacks across quantized neural networks with different architectures and bitwidths. As illustrated in Figure 1, our methodology consists of a series of controlled experiments designed to examine how factors like quantization level, spatial placement, gradient alignment, and patch size contribute to the success of patch-based attacks in QNNs. Each experiment targets a specific hypothesis about the behavior of QNNs under adversarial conditions, focusing on factors that could affect the transferability and resilience of patch-based attacks.
@@ 3.1 Feature Map Analysis Across Bit Widths:
This experiment investigates whether patch-based attacks generate distinct, localized feature activations that persist across different bit widths. By analyzing these activations, we aim to understand why patches are resilient to quantization noise and if the model’s focus on the adversarial patch is preserved despite reduced precision. If patches indeed produce prominent activations consistently across bit widths, it would reveal a critical vulnerability in quantized models, suggesting that localized adversarial signals remain impactful even under heavy quantization.
Objective: To determine whether patches induce strong, localized feature representations that remain consistent across varying bit widths, thereby sustaining the model’s focus on the patch regardless of quantization.
Method: We evaluate models quantized to different bit widths (32-bit, 8-bit, 4-bit, 2-bit) using QAT. Clean and patched inputs are passed through these models, and feature maps are extracted from key intermediate layers. These maps are then visualized to identify differences in representation and the persistence of patch-induced features.
@@ 3.2 Effect of Quantization Levels on Patch Transferability:
Objective: To assess the impact of quantization levels on the effectiveness of patch-based attacks and determine whether quantization strength directly influences patch transferability.
Method: We generate patch-based attacks using a 32-bit model and then evaluate these patches on the same model quantized to various bit widths (e.g., 8-bit, 4-bit, 2-bit) using both QAT and PTQ. The attack success rate is measured at each bit width to observe any reduction in effectiveness as bit width decreases.
@@ 3.3 Spatial Sensitivity of Patches:
Objective: To determine whether the effectiveness of patch-based attacks relies on precise spatial positioning rather than solely on their visual pattern, thereby revealing any specific spatial vulnerabilities within the structure of quantized neural networks.
Method: We generate patches on a 32-bit model and evaluate their effectiveness on QAT models quantized to 8, 5, 4, and 2 bits. For each test, we apply slight spatial modifications to the patch, including shifting by a few pixels, rotating, and scaling. We then record the attack success rate for each modification to assess the impact of these spatial adjustments on model vulnerability.
@@
Objective: To determine whether gradients associated with effective patch-based attacks exhibit stronger alignment between full-precision and quantized models compared to gradients for pixel-level adversarial perturbations.
Method: We generate gradients for both patch-based and pixel-level attacks on a 32-bit model and then measure the cosine similarity between these gradients and those on 8-bit, 4-bit, and 2-bit QAT models.
@@ 3.5 Transferability of Patch Across Architectures:
Objective: To evaluate if patch-based attacks generated on one architecture maintain high effectiveness when transferred to models with different architectures and varying quantization levels. This will help determine if the adversarial effectiveness of patches depends on architecture-specific representations or if they exploit more generalizable features that make them effective across architectures.
Method: We generate adversarial patches on a base architecture (e.g., Resnet-20 at 32-bit) with high attack success. We then transfer these patches to models with different architectures (e.g., Resnet-56, VGG-16, VGG-19) quantized using QAT at bitwidths such as 8-bit, 5-bit, 4-bit, and 2-bit. We measure and record the attack success rate across architectures and bitwidths.
@@ 3.6 Impact of Patch Size and Visibility:
Objective: To test if larger patches have higher success rates across quantized models, suggesting that the impact of patches is due to their high perceptual visibility.
Method: We generate patch-based attacks with varying patch sizes and opacity levels on a 32-bit model. We evaluate the success rate of each patch size and visibility level across QAT and PTQ models with 8 and 4-bit quantization.
@@ 4.1 Experimental Setup
Dataset: We use the CIFAR-10 dataset [19], which comprises 60,000 images evenly distributed across 10 classes.
Model Architectures: Our experiments cover a range of architectures including ResNet-20 [11], ResNet-56 [11], VGG-16 [31], VGG-19 [31], and AlexNet [20].
Quantization Techniques: We implement two quantization methods—Quantization-Aware Training and Post-Training Quantization—to examine model robustness at varied bitwidths (32-bit, 8-bit, 5-bit, 4-bit, and 2-bit).
Training Details: All models were trained with a batch size of 128 using Stochastic Gradient Descent (SGD) with a momentum of 0.9 and weight decay set to $1\times 10^{-4}$. Each model was trained for 50 epochs to ensure adequate convergence across quantization levels and architectures.
Patch-based Attacks: We employ two widely recognized patch-based attack techniques: LAVAN [17] and Adversarial Patch (GoogleAP) [4]. Further details are presented in the supplementary material.
@@ 4.2 Experiment 1: Feature Map Analysis Across Bit Widths
The feature maps in Figure 2 for patched inputs reveal distinct, high-activation regions at the patch location across all bitwidths. In each examined layer, the presence of the patch generates prominent localized activations, which remain visible even at the lowest bitwidths. This observation suggests that the patch induces a strong, persistent signal that is only partially attenuated by quantization.
Although these high-activation regions are consistent in patched inputs, the intensity of activations diminishes as the bitwidth decreases, particularly in deeper layers such as Conv2 and Conv3. In the 2-bit model, the activations appear more diffuse and less sharply defined. This reduction in intensity indicates that quantization affects the precision of the model’s response to the patch but does not fully neutralize the adversarial influence of the patch.
@@
ResNet-20 85.45 83.23 81.29 76.96 72.19
ResNet-56 86.55 78.47 75.81 84.53 72.52
VGG-16 80.95 84.80 81.54 77.33 83.03
VGG-19 88.17 76.29 77.59 88.20 90.95
AlexNet 79.204 81.391 82.560 83.447 89.011
Table 1: Attack Success Rates (%) for LAVAN untargeted attacks (6x6 patch) across different QNNs on CIFAR-10.
The persistence of patch-induced activations across all bitwidths indicates that the adversarial signal generated by the patch is resilient to quantization. This robustness is due to the high-visibility patterns created by the patch, which remain effective even at lower precision levels. These findings are consistent with the next experiment results showing high attack success rates for patch-based attacks across varying quantization levels, suggesting that the patch leverages strong, localized feature disruptions that quantization alone cannot effectively neutralize.
As shown in Table 1, attack success rates remain consistently high across all bitwidths, with mean success rates exceeding 70% even at the lowest bitwidth (2-bit). This resilience across quantization levels suggests that patch-based attacks maintain their adversarial impact despite significant precision reduction, indicating that the patch’s effect is robust to quantization.
Although attack success rates do decline slightly as bitwidth decreases, the reduction is generally modest. For example, ResNet-20 shows a decrease from 85.45% at 32-bit to 72.19% at 2-bit, while ResNet-56 drops from 86.55% at 32-bit to 72.52% at 2-bit. This pattern suggests that while quantization weakens the patch’s influence to some extent, it is insufficient to neutralize the attack entirely. The patch’s high visibility and distinct features continue to drive effective adversarial responses even under low-precision constraints.
The high success rates across bitwidths indicate that quantization alone provides limited defense against patch-based attacks, particularly for visible, localized perturbations like those in LAVAN attacks. The minimal drop in effectiveness suggests that quantized models remain vulnerable to such adversarial patterns.
4.4 Experiment 3: Spatial Sensitivity of Patches
32-bit 8-bit 5-bit 4-bit 2-bit
Original 88.17 76.29 77.59 88.2 90.95
Shifted (2,2) 89.241 75.385 77.844 82.609 89.041
Shifted (8,8) 87.342 82.564 80.24 79.348 89.954
Shifted (18,18) 81.646 81.538 79.042 81.522 79.452
Table 2: Attack Success Rates (%) for VGG-19 on CIFAR-10 across different bitwidths (32, 8, 5, 4, and 2-bit) at various patch positions.
As shown in Table 2, the original patch maintains high success rates across all bitwidths, from 88.17% on the 32-bit model to 90.95% on the 2-bit model. Minor shifts in patch position, such as (2,2) and (8,8), do not significantly impact attack effectiveness, with success rates remaining comparable to those of the original patch across quantized versions. For instance, the (2,2) shifted patch achieves an 89.24% success rate at 32-bit, and this robustness persists across lower bitwidths, indicating resilience to small spatial adjustments.
When the patch is shifted by a more substantial margin (18,18), the attack success rate experiences a slight decrease, particularly in the 32-bit model (81.65%) and the 2-bit model (79.45%). Despite this reduction, the success rates remain relatively high across all bitwidths, suggesting that the patch’s adversarial effectiveness is only marginally impacted by larger positional changes.
We also tested other spatial modifications, including rotation and scaling, and observed similar trends in attack robustness. Detailed results for these variations and other tested architectures are provided in the supplementary material. The results indicate that patch-based attacks are highly robust to spatial perturbations, showing only minor reductions in success rates even with significant shifts in patch position. This supports the hypothesis that patch attacks rely on high-visibility, localized patterns rather than precise spatial alignment.
@@ 4.4 Experiment 3: Spatial Sensitivity of Patches
4.6 Experiment 5: Transferability of Patch Across Architectures
As shown in Table 3, attack success rates remain consistently high across different architectures, even when patches are transferred between models. For example, a patch generated on ResNet-20 achieves an 84.03% success rate on 32-bit ResNet-56 and 76.33% on 32-bit VGG-19. Similarly, patches created on VGG-19 and VGG-16 maintain high success rates when tested on ResNet architectures, indicating strong cross-architecture transferability of patch-based attacks. This suggests that patch-based attacks are not heavily reliant on architecture-specific features, likely due to their highly visible and localized impact.
Quantization appears to have only a minor impact on the effectiveness of transferred patches. For instance, when transferring a patch from ResNet-20 to ResNet-56, success rates remain high even at 2-bit (68.42%). Similarly, transferring a patch from ResNet-56 to VGG-19 yields a high success rate of 88.89% at 2-bit. This minimal drop in effectiveness across bitwidths suggests that quantization does not significantly hinder the transferability of patch-based attacks, highlighting the resilience of these attacks to precision reduction. Additional results are provided in the supplementary material.
Resnet-56 VGG-19
32-bit 8-bit 5-bit 4-bit 2-bit 32-bit 8-bit 5-bit 4-bit 2-bit
Resnet-20 84.03 77.08 77.83 81.39 68.42 76.33 75.77 76.44 80.34 90.53
Resnet-20 VGG-19
32-bit 8-bit 5-bit 4-bit 2-bit 32-bit 8-bit 5-bit 4-bit 2-bit
Resnet-56 84.11 89.03 77 81.99 72.42 78.7 76.29 75.29 83.15 88.89
Resnet-20 VGG-16
32-bit 8-bit 5-bit 4-bit 2-bit 32-bit 8-bit 5-bit 4-bit 2-bit
VGG-19 85.43 86.45 74.96 84.44 72.79 71.04 87.35 85.87 76.22 85.31
VGG-19 Resnet-56
32-bit 8-bit 5-bit 4-bit 2-bit 32-bit 8-bit 5-bit 4-bit 2-bit
VGG-16 77.64 90.14 82.82 84.38 73.38 80.71 80.32 73.94 85.45 90.20
Table 3: Attack Success Rates (%) transfer across QNNs with different bitwidths (32, 8, 5, 4, 2-bit) and architectures (Resnet-20, Resnet-56, VGG-19 and VGG-16) on CIFAR-10.
The consistently high attack success rates across different architectures and bitwidths suggest that patch-based attacks exploit generalized, high-visibility features that are not strongly dependent on specific architectural representations. This characteristic makes them highly transferable and effective across a range of model architectures. These findings highlight a systemic vulnerability in quantized neural networks to patch-based attacks, as their success does not rely on the particular internal feature structures of individual architectures.

### v2

@@ Abstract
Quantized neural networks (QNNs) are increasingly used for efficient deployment of deep learning models on resource-constrained platforms, such as mobile devices and edge computing systems. While quantization reduces model size and computational demands, its impact on adversarial robustness—especially against patch-based attacks—remains inadequately addressed. Patch-based attacks, characterized by localized, high-visibility perturbations, pose significant security risks due to their transferability and resilience. In this study, we systematically evaluate the vulnerability of QNNs to patch-based adversarial attacks across various quantization levels and architectures, focusing on factors that contribute to the robustness of these attacks. Contrary to the expectation that quantization might enhance adversarial defenses, our results show that QNNs remain highly susceptible to patch attacks due to the persistence of distinct, localized features within quantized representations. To address this vulnerability, we propose Quantization-Aware Defense Training (QADT), a novel defense strategy that enhances QNN resilience by integrating adversarial patch augmentation and simulated quantized conditions during training. Our results demonstrate that QADT significantly improves the robustness of quantized models against patch-based attacks, reducing attack success rates by a substantial margin. These findings underscore the need for quantization-aware defenses that address the specific challenges posed by patch-based attacks and contribute to a deeper understanding of adversarial robustness in QNNs, guiding future research in developing secure, quantization-compatible defenses for real-world applications.
@@ 1 Introduction
As deep learning models are increasingly deployed in resource-constrained environments, such as mobile devices and edge computing platforms, quantization has emerged as a critical method to reduce model size and computational load. By lowering the bitwidth of the weights and activations, quantization enables models to operate more efficiently, preserving accuracy while significantly reducing resource demands [29]. QNNs are now central to applications ranging from autonomous systems [25, 18] to healthcare [37] and Internet of Things (IoT) devices [34, 12, 16]. However, as QNNs gain prominence, ensuring their robustness against adversarial attacks, a long-standing vulnerability in deep learning, remains an essential concern for secure deployment.
Among adversarial threats, patch-based attacks pose a unique and formidable challenge [9]. Unlike conventional adversarial perturbations that rely on subtle, dispersed changes, patch attacks introduce highly visible, localized perturbations that disrupt model predictions by “hijacking” the model’s focus. These attacks are especially problematic because of their transferability and resilience across diverse configurations, making them highly effective in physically realizable real-world scenarios. The distinct nature of patch-based attacks raises pressing questions about the robustness of QNNs against these localized, high-intensity adversarial patterns.
Quantization is often viewed as a potential defense against adversarial attacks, with evidence suggesting that reduced precision may mitigate some types of adversarial perturbations [2, 31, 7]. However, the effectiveness of quantization in countering patch-based attacks remains largely untested. Given that patch attacks rely on distinct and highly visible features to exploit model vulnerabilities, it is reasonable to suspect that they may retain their efficacy even in quantized models. Most existing studies on adversarial robustness in QNNs have focused on dispersed or pixel-level attacks, overlooking the unique challenges presented by patch-based attacks.
@@
  • We systematically evaluate how patch-based attacks affect feature representations at different quantization levels. Our findings reveal that quantization fails to effectively disrupt the adversarial signal of patches, as their high-visibility features persist across bit widths. This persistence underscores a critical vulnerability in QNNs.
@@
  • Based on our findings, we discuss key strategies to enhance the robustness of QNNs against patch-based adversarial attacks and propose Quantization-Aware Defense Training (QADT), a defense to increase QNN resilience to patch-based attacks.
@@ 2 Related Work
QNNs are widely adopted to reduce model size and computational costs, making them suitable for mobile devices and edge applications [29]. Quantization-Aware Training (QAT) [13, 30, 38, 23] and Post-Training Quantization (PTQ) [28, 24, 14, 35] are two prominent approaches for implementing low-bitwidth models, each introducing varying degrees of robustness against adversarial attacks [2, 31, 7]. Recent studies indicate that while quantization may disrupt the gradient landscape, creating potential resistance to some pixel-level perturbations [8, 27, 4], its effectiveness against more structured adversarial patterns remains uncertain [37]. Research on QNNs has focused mainly on dispersed, pixel-level adversarial attacks [21, 36], leaving a gap in understanding how QNNs respond to localized, high-visibility attacks like patch-based adversarial patterns.
@@ 3 Methodology
In this study, we systematically investigate the robustness and transferability of patch-based adversarial attacks across quantized neural networks with different architectures and bitwidths. As illustrated in Figure 1, our methodology consists of a series of controlled experiments designed to examine how factors such as quantization level, spatial placement, gradient alignment, and patch size contribute to the success of patch-based attacks in QNNs.
@@ 3.1 Feature Map Analysis Across Bit Widths:
This experiment investigates whether patch-based attacks generate distinct localized feature activations that persist across different bit widths. By analyzing these activations, we aim to understand why patches are resilient to quantization noise and if the model’s focus on the adversarial patch is preserved despite reduced precision. If patches indeed produce prominent activations consistently across bit widths, it would reveal a critical vulnerability in quantized models, suggesting that localized adversarial signals remain impactful even under heavy quantization.
@@ 4.1 Experimental Setup
Dataset: We use the CIFAR-10 dataset [19] and the ImageNet dataset [20].
Model Architectures: Our experiments cover a range of architectures including ResNet56, 34, 20, 18 [11], VGG-19, 16 [32], AlexNet [20], Inception-v3 [33], and the vision transformer Swin-S [26].
Quantization Techniques: We use different quantization methods: Quantization-Aware Training and Post-Training Quantization to examine model robustness at varied bitwidths (32-bit, 8-bit, 5-bit, 4-bit, and 2-bit). We also use different calibration techniques such as MinMax and Percentile.
Patch-based Attacks: We employ three widely recognized patch-based attack techniques: LAVAN [17], Adversarial Patch GoogleAP (GAP) [3] and Deformable Patch Representation (DPR) attack [5]. Further details are presented in the supplementary material.
@@ 4.2 Experiment 1: Feature Map Analysis Across Bit Widths
We evaluate models quantized to different bit widths (32-bit, 8-bit, 4-bit, 2-bit) using QAT. Clean and patched inputs are passed through these models and feature maps are extracted from key intermediate layers. These maps are then visualized to identify differences in representation and persistence of patch-induced features.
@@
The feature maps in Figure 2 for patched inputs reveal distinct high-activation regions at the patch location across all bitwidths. In each examined layer, the presence of the patch generates prominent localized activations, which remain visible even at the lowest bitwidths. This observation suggests that the patch induces a strong, persistent signal that is only partially attenuated by quantization.
The persistence of patch-induced activations across all bitwidths indicates that the adversarial signal generated by the patch is resilient to quantization. This robustness is due to the high-visibility patterns created by the patch, which remain effective even at lower precision levels. These findings are consistent with the results of the next experiment showing high attack success rates for patch-based attacks across varying quantization levels, suggesting that the patch leverages strong, localized feature disruptions that quantization alone cannot effectively neutralize.
In this experiment, we generate patch-based attacks using a 32-bit model and then evaluate these patches on the same model quantized to various bit widths (e.g., 8-bit, 4-bit, 2-bit) using both QAT and PTQ. The attack success rate is measured at each bit width to observe any reduction in effectiveness as the bit width decreases.
As shown in Table 1, attack success rates remain consistently high across all bitwidths, with mean success rates exceeding 70% even at the lowest bitwidth (2-bit). This resilience across quantization levels suggests that patch-based attacks maintain their adversarial impact despite a significant reduction in precision, indicating that the patch effect is robust to quantization.
Model 32-bit 8-bit 5-bit 4-bit 2-bit
ResNet-20 87.22 83.73 80.65 77.30 74.18
ResNet-56 86.43 83.24 79.94 76.22 73.08
VGG-19 88.95 85.56 82.28 79.81 77.19
VGG-16 87.17 84.73 81.45 78.29 76.67
AlexNet 82.92 79.72 77.31 75.88 73.54
Table 1: ASR (%) of LAVAN attack (6x6 patch) across different QNNs on CIFAR-10.
Although attack success rates do decline slightly as bitwidth decreases, the reduction is generally modest. For example, ResNet-20 shows a decrease from 87.22% at 32-bit to 74.18% at 2-bit. This pattern suggests that while quantization weakens the patch’s influence to some extent, it is insufficient to neutralize the attack entirely. The high visibility of the patch and distinct features continue to drive effective adversarial responses even under low-precision constraints.
We also evaluate the transferability of patch-based attacks on dynamically quantized models (8-bit) (Table 2). The results indicate that adversarial patches remain highly effective, confirming that vulnerabilities persist even under dynamic quantization.
Model ResNet-56 ResNet-20 VGG-19 VGG-16
LAVAN 84.03 83.29 76.33 71.58
GAP 82.40 53.76 54.12 41.78
Table 2: ASR (%) of LAVAN and GAP attacks (6x6 patches) across dynamically quantized models (8-bit) on CIFAR-10.
ResNet34 (QAT) ResNet18 (QAT)
NP 32-bit 5-bit 4-bit 2-bit 32-bit 5-bit 4-bit 2-bit
0.1 99.31 66.32 63.56 56.31 99.98 72.63 67.89 65.76
0.08 98.08 64.91 59.97 52.25 99.93 66.37 61.11 55.42
0.06 97.12 64.79 57.31 50.43 96.01 58.20 53.59 51.84
Table 3: ASR (%) of LAVAN attack with different noise percentages (NP) across different QNNs on ImageNet.
Model Quantization 32-bit 8-bit 4-bit 2-bit
Inception v3 QAT 89.10 60.12 55.32 50.66
Swin-S MinMax 91.80 62.11 - -
Swin-S Percentile 93.10 63.72 - -
Table 4: ASR (%) of LAVAN attack with a noise percentage 0.08 on ImageNet.
We extended our experiments to ImageNet using ResNet34, ResNet18, Inception-v3, all trained with QAT, along with Swin-S models quantized using PTQ with MinMax and Percentile calibration [6]. Quantization fails to efficiently disrupt patch-based attacks (Tables 3 and 4 )
We evaluated another patch-based attack, the Deformable Patch Representation (DPR) attack [5]. As shown in Table 5, DPR remains highly effective even on quantized ResNet34, achieving a 49.87% ASR at 2-bit.
Attack 32-bit 8-bit 4-bit 2-bit
DPR 83.46 59.32 55.82 49.87
Table 5: ASR (%) of DPR attack against ResNet34 on ImageNet.
4.4 Experiment 3: Spatial Sensitivity of Patches
We generate patches on a 32-bit model and evaluate their effectiveness on QAT models quantized to 8, 5, 4, and 2 bits. For each test, we apply slight spatial modifications to the patch, including shifting by a few pixels, rotating, and scaling. We then record the attack success rate for each modification to assess the impact of these spatial adjustments on model vulnerability.
@@ 4.4 Experiment 3: Spatial Sensitivity of Patches
Original 86.43 83.24 79.94 76.22 73.08
Shifted (2,2) 87.73 82.35 79.55 77.39 72.93
Shifted (8,8) 85.13 82.56 80.24 78.11 72.88
Shifted (18,18) 82.47 80.77 79.04 74.78 70.10
Rotated $5^{\circ}$ 86.39 82.98 80.81 76.16 72.84
Rotated $10^{\circ}$ 86.24 82.68 80.45 76.79 72.20
Rotated $20^{\circ}$ 84.38 81.50 79.85 75.17 71.73
Rotated $30^{\circ}$ 82.98 80.71 78.19 74.82 70.39
Table 6: ASR (%) of LAVAN attack against ResNet-56 (full precision and quantized) on CIFAR-10 at various patch positions and rotation angle.
As shown in Table 6, the original patch maintains high success rates across all bitwidths, from 86.43% on the 32-bit model to 73.08% on the 2-bit model. Minor shifts in patch position, such as (2,2) and (8,8), do not significantly impact attack effectiveness, with success rates remaining comparable to those of the original patch across quantized versions.
Patches maintain high attack success rates across all tested rotation angles, demonstrating resilience to spatial orientation changes. This robustness persists across different quantization levels, although success rates decline slightly at lower bit widths (e.g., 2-bit).
@@
We generate gradients for patch-based and pixel-level attacks on a 32-bit model and then measure the cosine similarity between these gradients and those on 8-bit, 4-bit, and 2-bit QAT models.
@@
4.6 Experiment 5: Transferability of Patches Across Architectures
We generate adversarial patches on a base architecture (e.g., Resnet-20 at 32-bit) with high attack success. We then transfer these patches to models with different architectures (e.g., Resnet-56, VGG-16, VGG-19) quantized using QAT at bitwidths such as 8-bit, 5-bit, 4-bit, and 2-bit. We measure and record the attack success rate across architectures and bitwidths.
As shown in Table 7, attack success rates remain consistently high across different architectures, even when patches are transferred between models. For example, a patch generated on ResNet-20 achieves an 84.17% success rate on 32-bit ResNet-56 and 78.82% on 32-bit VGG-19. Similarly, patches created on VGG-19 and VGG-16 maintain high success rates when tested on ResNet architectures, indicating strong cross-architecture transferability of patch-based attacks. This suggests that patch-based attacks are not heavily reliant on architecture-specific features, likely due to their highly visible and localized impact.
Quantization appears to have only a minor impact on the effectiveness of the transferred patches. For instance, when transferring a patch from ResNet-20 to ResNet-56, success rates remain high even at 2-bit (75.21%). Similarly, transferring a patch from ResNet-56 to VGG-19 yields a high success rate of 71.22% at 2-bit. This minimal drop in effectiveness across bitwidths suggests that quantization does not significantly hinder the transferability of patch-based attacks, highlighting the resilience of these attacks to precision reduction. Additional results are provided in the supplementary material.
Res.56 VGG-19
32-bit 8-bit 4-bit 2-bit 32-bit 8-bit 4-bit 2-bit
Res.20 84.17 79.62 77.66 75.21 78.82 74.53 72.15 70.21
Res.20 VGG-19
32-bit 8-bit 4-bit 2-bit 32-bit 8-bit 4-bit 2-bit
Res.56 84.11 77.67 75.33 71.76 77.43 75.09 73.82 71.22
Res.20 VGG-16
32-bit 8-bit 4-bit 2-bit 32-bit 8-bit 4-bit 2-bit
VGG19 83.23 80.87 78.11 75.32 85.32 80.42 78.23 76.44
VGG-19 Res.56
32-bit 8-bit 4-bit 2-bit 32-bit 8-bit 4-bit 2-bit
VGG16 83.29 78.48 76.65 74.39 80.55 78.93 75.34 73.87
Table 7: ASR (%) transfer across different QNNs with different bitwidths and architectures on CIFAR-10.